Person got it wrong the first time I posted this.

Iconic memory is a type of memory that holds visual information for about half a second (0.5 seconds). To demonstrate this type of memory, participants were shown three rows of four letters for 50 milliseconds. They were then asked to recall as many letters as possible, with a 0-, 0.5-, or 1.0-second delay before responding. Researchers hypothesized that longer delays would result in poorer recall. The number of letters correctly recalled is given in the table.

Delay Before Recall
0 0.5 1
10 10 2
11 8 2
8 4 7
7 3 3
12 5 5
6 6 5

(a) Complete the F-table. (Round your values for MS and F to two decimal places.)

Source of Variation SS df MS F
Between groups x x x x
Within groups (error) x x x
Total x x


(b) Compute Tukey's HSD post hoc test and interpret the results. (Assume alpha equal to 0.05. Round your answer to two decimal places.)

The critical value is ___ for each pairwise comparison.


Which of the comparisons had significant differences? (Select all that apply.)

The null hypothesis of no difference should be retained because none of the pairwise comparisons demonstrate a significant difference.

Recall following no delay was significantly different from recall following a one second delay.

Recall following a half second delay was significantly different from recall following a one second delay.

Recall following no delay was significantly different from recall following a half second delay.
Answer #1

using excel addins data analysis tool for one way anova,output obtained is

a)

ANOVA: Single Factor
SUMMARY
Groups Count Sum Average Variance
0 6 54 9 5.6000
0.5 6 36 6 6.8000
1 6 24 4 4.0000
ANOVA
Source of Variation SS df MS F P-value F crit
Between Groups 76.0 2 38.00 6.95 0.01 3.68
Within Groups 82.0 15 5.47
Total 158.0 17
Level of significance 0.05

b)

critical value = q0.05,3,15 * √(MSW/n) = 3.6734*√(5.467/6) = 3.51

if absolute difference of means>critical value, difference is significant,otherwise not

c)

Tukey-Kramer Multiple Comparisons
Sample Sample Absolute Std. Error Critical
Group Mean Size Comparison Difference of Difference Range Results
group 1= 0 9 6 Group 1 to Group 2 3 0.954521404 3.5063 Means are not different
group2=0.5 6 6 Group 1 to Group 3 5 0.954521404 3.5063 Means are different
group3=1 4 6 Group 2 to Group 3 2 0.954521404 3.5063 Means are not different
Other Data
Level of significance 0.05
Numerator d.f. 3
Denominator d.f. 15
MSW 5.466667
Q Statistic 3.6734

Recall following no delay was significantly different from recall following a one second delay. [ answer]